### v2

# Poor man’s transcendence for Frobenius traces of elliptic curves

Abstract.
Let $E$ be an elliptic curve without complex multiplication defined over $\mathbb{Q}$. Viewing the sequence of its Frobenius traces $(a_{p}(E))_{p}$ indexed by primes $p$ as an element in the ‘poor man’s adèle ring’, we prove its transcendence over $\mathbb{Q}$.
Key words and phrases:
Transcendence, poor man’s adèle ring, elliptic curve, Frobenius endomorphism
2020 Mathematics Subject Classification:
11J81 (primary), 11A41, 11J72, 11B83, 11G07, 13A35, 16U10 (secondary).
Many reincarnations of multiple zeta values are known to this date; one of them was introduced recently by Kaneko and Zagier in [6] (see also [4, 5]). This remarkable and innocently looking version called finite multiple zeta values lives in the universe of ‘poor man’s adèle ring’ [7]
$\mathcal{A}=\bigg{(}\prod_{p}(\mathbb{Z}/p\mathbb{Z})\bigg{)}\bigg{/}\bigg{(}\bigoplus_{p}(\mathbb{Z}/p\mathbb{Z})\bigg{)}.$
The elements of the ring are infinite vectors $\boldsymbol{t}=(t_{2},t_{3},t_{5},t_{7},\dots)=(t_{p})_{p\in\mathcal{P}}$ indexed by prime numbers, with each $t_{p}$ viewed as a residue in $\mathbb{Z}/p\mathbb{Z}$. The equivalence relation $\boldsymbol{t}\sim\boldsymbol{s}$ corresponding to $t_{p}\equiv s_{p}\bmod p$ for all sufficiently large primes $p$ (in particular, both $t_{p}\bmod p$ and $s_{p}\bmod p$ are well defined for such primes $p$); a few first components of $\boldsymbol{t},\boldsymbol{s}$ can remain undefined. The field $\mathbb{Q}$ of rational numbers is naturally embedded in $\mathcal{A}$ diagonally, so that $\mathcal{A}$ is regarded as a $\mathbb{Q}$-algebra via this embedding.
Some very basic open questions about finite multiple zeta values include the nonvanishing of the latter, which in turn motivate investigating irrationality and even transcendence in the ring $\mathcal{A}$. First examples of $\mathcal{A}$-transcendental numbers are discussed in [1, 8]. Though there are two competing concepts of algebraicity in $\mathcal{A}$ — a naïve (‘combinatorial’) one from [1] and a truly algebraic one introduced by Rosen in [9] (see also [10]), we focus our attention only on the latter. Roughly speaking, an algebraic element $\boldsymbol{t}$ in $\mathcal{A}$ originates from a $C$-finite sequence $(t_{m})_{m\gg 1}\in\mathbb{Q}^{\infty}$, that is, from a sequence satisfying a linear recursion with constant rational coefficients. More formally, we have the following criterion to detect the algebraicity.
Proposition 1 ([9, Theorem 1.1]).
Let $\boldsymbol{t}=(t_{p})_{p\in\mathcal{P}}\in\mathcal{A}$. The following conditions are equivalent.
  • (i)
    The element $\boldsymbol{t}$ is a finite algebraic number.
  • (ii)
    There exists a Galois extension $L/\mathbb{Q}$ and a map $\phi\colon\operatorname{Gal}(L/\mathbb{Q})\to L$ satisfying
    $\phi(\sigma\tau\sigma^{-1})=\sigma(\phi(\tau))\quad\text{for all}\;\sigma,\tau\in\operatorname{Gal}(L/\mathbb{Q})$
    such that
    $(t_{p})_{p\in\mathcal{P}}=(\phi(\operatorname{Frob}_{p})\bmod p)_{p\in\mathcal{P}},$
    where $\operatorname{Frob}_{p}$ is the Frobenius map of $L$ at the prime $p$.
As we have already indicated in [8], there are somewhat natural elements in $\mathcal{A}$ for which the transcendence is a consequence of existing difficult conjectures. One such a family of potentially transcendental examples corresponds to the sequence of Frobenius traces $\boldsymbol{a}=(a_{p})_{p\in\mathcal{P}}$ attached to an elliptic curve over $\mathbb{Q}$; we demonstrate the $\mathcal{A}$-irrationality of these $\boldsymbol{a}$ in [8]. The principal goal of this note is establish their $\mathcal{A}$-transcendence unconditionally in the case of elliptic curve without complex multiplication.
Theorem 1.
Let $E$ be an elliptic curve defined over $\mathbb{Q}$, without complex multiplication. Then the sequence of its Frobenius traces $\boldsymbol{a}=(a_{p}(E))_{p\in\mathcal{P}}$ is $\mathcal{A}$-transcendental.
Proof.
Assume for a contradiction that $\boldsymbol{a}$ is $\mathcal{A}$-algebraic. By part (ii) of Proposition 1 there exist data $L/\mathbb{Q}$ and $\phi\colon\operatorname{Gal}(L/\mathbb{Q})\to L$; collect the finite set $\{b_{1},\dots,b_{k}\}$ of algebraic numbers given by the image of $\operatorname{Gal}(L/\mathbb{Q})$ in $L$. The congruence
$a_{p}\equiv b_{i}\bmod p\quad\text{for some}\;i\in\{1,\dots,k\}$
means that for all $p$ except finitely many, the prime $p$ divides the norm $\mathbb{N}(a_{p}-b_{i})$ for some $i\in\{1,\dots,k\}$.
Let $E_{\ell}$ denote the set of $\overline{\mathbb{Q}}$-points on the elliptic curve $E$ of order $\ell$. The action of Galois group $\operatorname{Gal}(L/\mathbb{Q})$ on $E_{\ell}$ gives a representation $\rho_{\ell}\colon\operatorname{Gal}(L/\mathbb{Q})\to\operatorname{Aut}(E_{\ell})\cong GL_{2}(\mathbb{Z}/\ell\mathbb{Z})$. Since $E$ has no complex multiplication, Serre’s open mapping theorem [11] guarantees that, for all but finitely many primes $\ell$, the image is the whole of $GL_{2}(\mathbb{Z}/\ell\mathbb{Z})=GL_{2}(\mathbb{F}_{\ell})$. We fix any prime $\ell$ of this quality subject to the additional condition that it is relatively prime to all integers in the collection $\{b_{1},\dots,b_{k}\}$.
Now choose $X$ to be large and eliminate all primes $p\leq X$ for which $a_{p}=0$; there are just $O(X^{3/4})$ of them, as a consequence of Kaneko’s result from [3] (see [2] for the details of this deduction). Then keep only those remaining primes for which $a_{p}\equiv 0\bmod\ell$; the latter condition means that the Frobenius of $L$ at $p$ maps to a matrix in $GL_{2}(\mathbb{F}_{\ell})$ with the zero trace — the proportion of such primes is exactly the proportion of zero-trace matrices, hence it is positive (in other words, it forms a Chebotarev set):
#$\#\{p\leq X:a_{p}\neq 0,\,a_{p}\equiv 0\bmod\ell\}\gg X/\log X.$
Let $i$ be such that $b_{i}$ is nonzero and integer. For our primes the equality $a_{p}=b_{i}$ cannot happen, because $a_{p}\equiv 0\bmod\ell$ and $(\ell,b_{i})=1$ by the choice of $\ell$. Thus, the relation
$p\mid\mathbb{N}(a_{p}-b_{i})\;\text{for some}\;i\in\{1,\dots,k\}\;\text{such that}\;b_{i}\neq 0$
is a genuine divisibility relation in that the norm $\mathbb{N}(a_{p}-b_{i})$ is not zero. Since $p\leq X$, we have $|a_{p}|\leq 2\sqrt{X}$, so that there are $O(X^{1/2})$ ways of choosing $a_{p}$. Each time this choice is performed, the nonzero integer $\mathbb{N}(a_{p}-b_{i})$ has size $X^{O(1)}$, hence it has at most $O(\log X)$ prime factors. Summing up over all $a_{p}$ and all possible $i$, the number of primes coming out of these divisibilities is at most
$O(\sqrt{X}\log X).$
At the same time $\gg X/\log X$ such primes dividing the norms are involved. The comparison of the sizes gives a contradiction for large $X$. ∎
Finally, we would like to notice that the $\mathcal{A}$-transcendental examples in Theorem 1 are very different in nature from the $q$-Fibonacci ones treated in [1, 8]. However what tended to work in both situations is rough counts on how often the entries repeat on a scale of up to $X$ versus how many primes are available. This is what has achieved progress. One can expect that the argument generalises to other situations. The elements in $\mathcal{A}$ attached to Frobenius traces suggest considering the sequences $(f(p))_{p\in\mathcal{P}}$ for multiplicative functions $f\colon\mathbb{Z}_{>0}\to\mathbb{Z}$. At the same time, many of those are $\mathcal{A}$-rational for trivial reasons — consider $f(n)=\sum_{d\mid n}d^{k}$ with $k\in\mathbb{Z}_{\geq 0}$ as such an example; the latter choice of $f(n)$ still expects to generate $\mathcal{A}$-transcendental elements $(f(p+c))_{p\in\mathcal{P}}$ via a shift by a fixed $c\in\mathbb{Z}\setminus\{0\}$, losing the multiplicativity though.
Acknowledgements.
We thank Masanobu Kaneko, Pieter Moree and Igor Shparlinski for their comments related to this note.
F.L. was partly supported by the 2024 ERC Synergy Project DynAMiCs. W.Z. acknowledges support of the Max Planck Institute for Mathematics (Bonn, Germany) during his stay in April–June 2025.